# Solidification estimates for random walks on supercritical percolation clusters

### v1

@@ 4.2. Resonance set
(recall $\mathbb{P}^{u}_{0}[~\cdot~]=\mathbb{P}^{u}[~\cdot\mid x\in\mathscr{S}_{\infty}]$). To this end, we let $(\ell_{n})_{n\geq 0}$, $(r_{n})_{n\geq 0}$, $(L_{n})_{n\geq 0}$ be defined as in [2, Theorem 6.4], and then apply it to prove the adjusted version of Lemmas 5.2 and 5.4 in [23], with $\frac{3}{4}\eta(u)$ and $\frac{5}{4}\eta(u)$ replaced by $\eta_{1}(\alpha)$ and $\eta_{2}(\alpha)$ respectively. The adjustment is valid since only the relation $\frac{3}{4}\eta(u)<\eta(u)<\frac{5}{4}\eta(u)$ has been used in the proofs to apply the ergodic theorem (see [23, (5.1)]). Hence there exists $C_{1}(\alpha)$ and $C_{2}(\alpha,\ell_{0}(\alpha))$ (the dependence on $u$ is omitted) such that for all $\ell_{0},r_{0}\geq C_{1}(\alpha)$, $L_{0}\geq C_{2}(\alpha,\ell_{0}(\alpha))$, and $n\geq 0$,
@@
It then follows by the above inequality, the definition of $\mathscr{H}_{K,s}^{\alpha}(x_{s})$, and S2, that

### v2

@@ 4.2. Resonance set
(recall $\mathbb{P}^{u}_{0}[\,\cdot\,]=\mathbb{P}^{u}[\,\cdot\,|\,{0}\in\mathscr{S}_{\infty}]$). To this end, we let $(\ell_{n})_{n\geq 0}$, $(r_{n})_{n\geq 0}$, $(L_{n})_{n\geq 0}$ be defined as in [2, Theorem 6.4], and then apply it to prove the adjusted version of Lemmas 5.2 and 5.4 in [23], with $\frac{3}{4}\eta(u)$ and $\frac{5}{4}\eta(u)$ replaced by $\eta_{1}(\alpha)$ and $\eta_{2}(\alpha)$ respectively. The adjustment is valid since only the relation $\frac{3}{4}\eta(u)<\eta(u)<\frac{5}{4}\eta(u)$ has been used in the proofs to apply the ergodic theorem (see [23, (5.1)]). Hence there exists $C_{1}(\alpha)$ and $C_{2}(\alpha,\ell_{0}(\alpha))$ (the dependence on $u$ is omitted) such that for all $\ell_{0},r_{0}\geq C_{1}(\alpha)$, $L_{0}\geq C_{2}(\alpha,\ell_{0}(\alpha))$, and $n\geq 0$,
@@
It then follows by the above inequality, the definition of $\mathscr{H}_{K,s}^{\alpha}(x_{s})$, and S1, that
@@
Acknowledgements While this work was written, AC was associated with INdAM (Istituto Nazionale di Alta Matematica “Francesco Severi”) and the group GNAMPA. MN was partially supported by Hong Kong RGC (Research Grants Council) grants ECS 26301824 and GRF 16303825. The authors wish to thank the referee for the thorough review of the article and for valuable suggestions.